# justification=centering

### v1

@@ Abstract
Methods. We generate detection maps and contrast curves for 219 observations in the F150 SHINE sample using the RSM algorithm. A clustering approach based on environmental parameters is used to group observations with similar noise characteristics. We propose two methods for defining radial detection thresholds in the RSM maps: fitting a log-normal distribution to the post-processed noise and maximizing the F1 score. We also assess the performance of various combinations of post-processing techniques within the RSM framework to identify optimal configurations.
Results. This study demonstrates the utility of clustering based on observational parameters, effectively distinguishing features like wind-driven halos and low-wind effects. Detection thresholds vary significantly across clusters, differing by up to a factor of 10, highlighting the importance of considering observational environments. Log-normal thresholds provide conservative, noise-aware limits, while F1 score-based thresholds offer observation-specific results, both showing compatibility overall. RSM improves detection limits by an average factor of two at $1\arcsec$ and five at inner working angles compared to standard PCA processing. This study reports more than 30 newly detected signals, including one promising candidate awaiting second-epoch confirmation.
@@ 1 Introduction
The field of high-contrast imaging (HCI) has undergone remarkable advancements in recent years, driven by innovations in coronagraphy, adaptive optics, sophisticated observing strategies, and advanced post-processing techniques. These developments have made direct imaging a powerful tool for detecting hot, massive Jupiter-like exoplanets at wide orbital separations, capable of capturing planets that are up to $10^{6}$ times dimmer than their host stars in the infrared. This capability has led to the detection and characterization of dozens of young exoplanets, offering invaluable insights into planetary formation processes (see 2023ASPC..534..799C, for a recent review). Pushing detection limits further to identify fainter, closer-in planets would bridge the gap between direct imaging and indirect methods, providing richer information about planetary formation pathways, enabling detailed atmospheric characterization of a larger sample of exoplanets, and expanding our understanding of exoplanet demographics.
@@
The Spectro-Polarimetric High-contrast Exoplanet Research (SPHERE, 2019A&A...631A.155B) instrument is a second-generation extreme adaptive optics system at the Very Large Telescope (VLT), equipped with advanced coronagraphs feeding imaging and spectroscopic cameras. SPHERE has been central to several exoplanet imaging campaigns, including the SPHERE High-contrast Imaging survey for Exoplanets (SHINE, 2017sf2a.conf..331C). SHINE targets around 400 young, nearby, and relatively bright stars to detect and characterize giant exoplanets, contributing to our understanding of their architectures and formation (see 2021A&A...651A..70D, for further details). The survey utilizes both of SPHERE’s near-infrared imaging cameras – the dual band imager IRDIS (2008SPIE.7014E..3LD) and the integral field spectrograph IFS (2010SPIE.7735E..0VC) –, employing angular differential imaging (ADI, 2006ApJ...641..556M) for high-contrast observations. The SHINE survey has enabled key discoveries, such as HIP 65426b (2017A&A...605L...9C) and PDS 70b (2018A&A...617L...2M), has refined constraints on giant planet occurrence rates (2021A&A...651A..71L), demographics of young exoplanets within 300 AU (2021A&A...651A..72V), and has spatially resolved several circumstellar disks (2016A&A...586L...8L; 2017A&A...601A...7F; 2016A&A...595A.114D). In their analysis of a 150-star subset from the SHINE survey, referred to as the F150 sample, 2021A&A...651A..71L evaluated the survey’s detection capabilities. Utilizing SPHERE’s IRDIS H-band mode, the median detection performance achieved $5\sigma$ contrasts of 11.8 mag at 200 mas and 13.1 mag at 800 mas. These results were obtained by applying standard post-processing techniques, including principal component analysis (PCA, Soummer_2012ApJ...755L..28S; 2012MNRAS.427..948A) and the template-based locally optimized combination of images (TLOCI, 2014SPIE.9148E..0UM).
@@
In recent years, more advanced post-processing algorithms such as ANDROMEDA (2015A&A...582A..89C), PACO (2018A&A...618A.138F; 2020A&A...637A...9F; 2024MNRAS.527.1534F), SODINN / NA-SODINN (2018A&A...613A..71G; 2023A&A...680A..86C), RSM (2020A&A...633A..95D; 2021A&A...646A..49D; 2021A&A...656A..54D), or 4S (2025AJ....169..194B) have been developed, with the goal of optimizing noise modeling and enhancing planetary signal detection – particularly for faint companions at close separations. Notably, 2023A&A...675A.205C applied PACO to the SHINE F150 sample, improving the median contrast limits by a factor five at inner working angles and a factor two overall, and demonstrating the impact of more refined statistical modeling.
@@
Building on this progress, we revisit the SHINE F150 sample using the RSM algorithm, which has shown excellent performance in the Exoplanet Imaging Data Challenge (EIDC, 2021arXiv210105080C) by achieving high F1 scores and a low false positive rate. Unlike spatial-only modeling approaches, RSM captures both the spatial and temporal evolution of pixel intensities across residual, derotated frames using a Markov regime-switching framework applied to outputs from multiple PSF subtraction techniques. This enables it to fully exploit the temporal structure of high-contrast imaging sequences, offering improved sensitivity and robustness in exoplanet detection. In this study, we define detection thresholds tailored to RSM maps, accounting for noise variations across observing conditions. We compare two thresholding strategies, one assuming a log-normal noise distribution, and another based on maximizing the F1 score, and assess their relative performance. We further evaluate how different post-processing algorithm combinations affect RSM sensitivity and present updated detection limits for the SHINE survey based on this reanalysis.
@@
The structure of this paper is as follows: Sect. 2 outlines the F150 sample selection criteria. Section 3 describes the observational conditions in the dataset. Section 4 explains the RSM algorithm and its application in this study. Section 5 examines two methods for setting detection thresholds and their atmospheric context. Section 6 presents contrast curves for different detection thresholds and post-processing combinations, highlighting the optimal approach within the RSM framework, and discusses the improvements in the SHINE detection limits. Section 7 presents the point sources detected by RSM in the F150 sample and compares them with those identified by other algorithms, such as PACO, including a discussion of newly detected candidates.
@@ 2 The data sample
In this study, we use the pre-reduced F150 sample from the SHINE survey, observed in H23 bands with ADI mode, covering an approximately $9\arcsec\times 9\arcsec$ field of view. The High Contrast Data Center (HC-DC, 2017sf2a.conf..347D; 2018A&A...615A..92G; 2019A&A...631A.155B; 2016SPIE.9908E..34M) pre-processed these observations, resulting in a dataset of 343 observations. Atmospheric conditions during these observations varied significantly, with seeing values ranging from $0\aas@@fstack{\prime\prime}4$ to $3\arcsec$ and Strehl ratios from 0.1 to 0.95. To ensure data quality and consistency, a pre-selection of datasets was necessary based on these atmospheric parameters.
The preliminary selection of data was based on several parameters, including seeing, Strehl ratio, raw contrast, number of frames, FWHM, and an assessment of the features in the PSF and the science cubes. Information like seeing, wind speed, coherence time were primarly derived from differential image motion monitor (DIMM, 1990A&A...227..294S) measurements, as recommended by 2017arXiv171005417M, while the Strehl ratio was empirically measured from the PSF. Additional parameters, such as raw contrasts for science cubes and frame quality assessments, were provided by the HC-DC reduction pipeline. In cases where DIMM seeing measurements were unavailable, they were inferred from the available seeing values provided by the SPHERE real-time computer SPARTA (2006SPIE.6272E..10F) using the relationship between DIMM and SPARTA seeing established in 2017arXiv171005417M.
To ensure a sample of stars with observing conditions ranging from fair to excellent, observations that met any of the following conditions were also discarded: seeing larger than $2\arcsec$, Strehl ratio less than 0.5, FWHM greater than 5.5 pixels, number of frames in the ADI sequence less than 40, and nonphysical raw contrasts at 500 mas (negative or greater than one). In case of multiple observations of the same star within five consecutive nights, only the observation with the best conditions was retained. Finally, we assessed the quality of both coronagraphic and non-coronagraphic cubes using a combination of tools: the HC-DC frame quality assessment, correlation functions from the VIP package (2017AJ....154....7G; 2023JOSS....8.4774C), and visual inspection. Cubes containing a majority of corrupted frames were discarded. In addition, science cubes in which the star was located outside the coronagraph mask, or where the off-axis PSF showed strong secondary lobes caused by the low-wind effect (see 2018SPIE10703E..2AM), were also excluded.
These filtering steps resulted in a total of 219 observations for 150 stars. These constitute the data sample used in this paper. The star names, along with the corresponding descriptive information and parameters, are all summarized in Table LABEL:tab:obs (see Appendix A).
@@ 3 Clustering based on environmental conditions
The quality of observations and specific features within coronagraphic images can be characterized by several environmental and instrumental parameters. Factors such as seeing, wind speed, coherence time, Strehl ratio, and raw contrast collectively indicate observation quality and provide insights into SPHERE-specific features. For instance, the wind-driven halo (2020A&A...638A..98C) arises under high wind velocity and short coherence times, while the low-wind effect (LWE, 2018SPIE10703E..2AM) appears at low wind speeds. These features can severely degrade SPHERE image quality, reducing the achievable contrast levels (see 2019Msngr.176...25C, for more details).
@@
We adopted the clustering method outlined in 2022A&A... [^6] A..33D, grouping observations with similar characteristics using the k-means clustering algorithm from the scikit-learn package. This approach not only helps identify distinct classes of SPHERE observations but also reduces computational load when optimizing PSF subtraction techniques and RSM parameters (as described in Section 4). The parameters selected for clustering include the number of frames in the ADI sequence, seeing, wind speed, parallactic angle range, Strehl ratio, raw contrast at 500 mas, and coherence time. Observations were grouped into six clusters. The number of clusters was determined by maximizing the silhouette score, a metric used to evaluate the quality of clustering. For each cluster, k-means automatically selected a representative observation, i.e., the dataset closest to the cluster centroid. In the following, the terms cluster center and representative observation are used interchangeably. The parameters for PSF subtraction and the RSM algorithm can then be optimized for these representative data sets, and generalized to other observations within the same cluster. To ensure unbiased parameter selection, we made sure that none of the cluster centers contained datasets with bright astrophysical signals. A pixel correlation test on science cubes validates the effectiveness of clustering in capturing noise features (see Appendix B), and supports the use of the k-means elected centers for parameter optimization. Figure 1 displays a $200\times 200$ pixel crop of representative individual frames from different clusters. These images illustrate the varying noise behavior across clusters, with some frames showing secondary lobes at close separations from the star indicative of the low-wind effect (Fig. 1a), while others exhibit strong elongated (aka butterfly) patterns characteristic of the wind driven halo (Fig. 1e).
@@
Figure 1: Image representation of single frames from various observations across different clusters, labeled (a) to (f) corresponding to clusters 1 to 6 presented in log scale.
@@
To visualize the clustering effectiveness based on selected parameters, we applied Principal Component Analysis (PCA) for dimensionality reduction (see Appendix C), enabling clear separation of clusters in the latent space. For a more physically grounded interpretation of the nature of the various clusters, Fig. 2 presents their visualization based on key environmental parameters, revealing distinct traits for each cluster:
@@
This successful clustering of data into meaningful categories will allow us to later analyze noise distributions and detection limits based on well-defined observational conditions, enabling a more targeted approach to interpreting results.
@@ 4 The RSM algorithm
The Regime Switching Model (RSM) is a well-established econometric algorithm that was successfully adapted to HCI by 2020A&A...633A..95D to enhance the detection of faint companions at small angular separations from their host stars. In the RSM algorithm, patches of pixels from derotated, residual frames after PSF subtraction are used to construct time series that capture spatial and temporal noise behavior at a given separation. When a planetary signal is present, these time series exhibit significant deviations from the noise patterns across the pixel patch and instead follow a planetary model that is both spatially and temporally structured. The algorithm models the data with two regimes: a noise regime, describing the mean and variance of noise at a given separation, and a planetary regime, incorporating the noise profile to which is added a model of the PSF whose intensity, $\beta$, is tuned. Each central pixel of the patch in a given annulus is assigned a likelihood of belonging to either regime. The RSM algorithm employs a Markov Chain model, where the probability of a pixel belonging to the planetary regime depends on the likelihood of the pixel-centered patch alignment with a planetary model, the probability of adjacent pixels, and a transition probability that connects these points. This integration of spatial and temporal constraints, combined with the transition probabilities, enables the algorithm to reliably distinguish speckles from true astrophysical signals. Additionally, the RSM algorithm utilizes outputs from various PSF subtraction techniques to construct a unified time series for each annulus. By drawing on the temporal noise diversity across different post-processing methods, this approach enhances the sensitivity to faint companions.
@@
RSM has demonstrated significant advancements over current post-processing techniques, as highlighted in multiple studies (2020A&A...633A..95D; 2021A&A...646A..49D; 2022A&A... [^6] A..33D). Furthermore, RSM shows excellent performance compared to other post-processing algorithms, achieving the lowest false positive rate in the Exoplanet Imaging Data Challenge (EIDC, 2021arXiv210105080C), underscoring its reliability in distinguishing planetary signals from noise.
@@
One of the most sensitive aspects of the RSM algorithm involves selecting optimal parameters for both PSF subtraction techniques and RSM itself, including the planetary flux multiplicative factor $\delta$ (where planetary flux is expressed as a multiple of the standard deviation of pixel intensity at a given separation, $\beta=\delta\sigma$, see equation 1 in 2020A&A...633A..95D), PSF model crop size, and the region used for noise estimation within an annulus. 2021A&A...656A..54D introduced an automated optimization process within RSM to identify these optimal settings by maximizing contrast for PSF subtraction parameters and minimizing false positives for RSM parameters using the reversed parallactic angles, thereby enhancing the algorithm’s reliability in exoplanet detection. In this study, we employ RSM with a simplified parameter set, as recommended by 2022A&A... [^6] A..33D. In the following, we refer to the value of each pixel in the final RSM map – previously defined as the RSM probability – as the RSM score, to avoid any potential ambiguity.
@@
To reduce computational load and mitigate potential issues with the RSM parameter optimization on certain pathological datasets, we propose to determine the optimal parameters for the cluster centers identified in Sect. 3 and to apply these settings consistently across all datasets within each cluster. Table 2 in Appendix D details the specific parameters for each cluster. To assess the reliability of this approach, Appendix E compares the cluster center parameters with those computed individually for each dataset, providing insights into their alignment and the robustness of the clustering methodology. We exclude from the optimization process the selection of the optimal combination of PSF subtraction techniques within the auto-RSM framework, as delivered by the opti-combination method described in 2021A&A...656A..54D. Instead, we keep this parameter adjustable and explore how varying combinations of PSF subtraction techniques impact noise distributions and contrast curves across separations, as discussed in Sect. 6.

### v2

@@ Abstract
Methods. We generated detection maps and contrast curves for 213 observations in the F150 SHINE sample using the RSM algorithm. A clustering approach based on environmental parameters was used to group observations with similar noise characteristics. We propose two methods for defining radial detection thresholds in the RSM maps: fitting a lognormal distribution to the post-processed noise and maximizing the F1 score. We also assessed the performance of various combinations of post-processing techniques within the RSM framework to identify optimal configurations.
Results. This study demonstrates the utility of clustering based on observational parameters, effectively distinguishing features such as wind-driven halos and low-wind effects. Detection thresholds vary significantly across clusters, differing by up to a factor of ten, highlighting the importance of considering observational environments. Lognormal thresholds provide conservative, noise-aware limits, while F1 score-based thresholds offer observation-specific results, with both showing compatibility overall. RSM improves detection limits by an average factor of two at $1\arcsec$ and five at inner working angles compared to standard principal component analysis processing. This study reports more than 30 newly detected signals, including one promising candidate awaiting second-epoch confirmation.
@@ 1 Introduction
The field of high-contrast imaging (HCI) has undergone remarkable advancements in recent years, driven by innovations in coronagraphs, adaptive optics, sophisticated observing strategies, and advanced post-processing techniques. These developments have made direct imaging a powerful tool for detecting hot, massive Jupiter-such as exoplanets at wide orbital separations, capable of capturing planets that are up to $10^{6}$ times dimmer than their host stars in the infrared. This capability has led to the detection and characterization of dozens of young exoplanets, offering invaluable insights into planetary formation processes (see Currie et al., 2023, for a recent review). Pushing detection limits further to identify fainter, closer-in planets would bridge the gap between direct imaging and indirect methods, providing richer information about planetary formation pathways, enabling detailed atmospheric characterization of a larger sample of exoplanets, and expanding our understanding of exoplanet demographics.
@@
The Spectro-Polarimetric High-contrast Exoplanet Research (SPHERE; Beuzit et al., 2019) instrument is a second-generation extreme adaptive optics system at the Very Large Telescope (VLT), equipped with advanced coronagraphs feeding imaging and spectroscopic cameras. SPHERE has been central to several exoplanet imaging campaigns, including the SPHERE High-contrast Imaging survey for Exoplanets (SHINE; Chauvin et al., 2017a). SHINE targets around 400 young, nearby, and relatively bright stars to detect and characterize giant exoplanets, contributing to our understanding of their architectures and formation (see Desidera et al., 2021, for further details). The survey utilizes both of SPHERE’s near-infrared imaging cameras – the dual band imager IRDIS (Dohlen et al., 2008) and the integral field spectrograph (IFS; Claudi et al., 2010) –, employing angular differential imaging (ADI; Marois et al., 2006) for high-contrast observations. The SHINE survey has enabled key discoveries, such as HIP 65426b (Chauvin et al., 2017b) and PDS 70b (Müller et al., 2018); has refined constraints on giant planet occurrence rates (Langlois et al., 2021) and demographics of young exoplanets within 300 AU (Vigan et al., 2021); and has spatially resolved several circumstellar disks (Lagrange et al., 2016; Feldt et al., 2017; de Boer et al., 2016). In their analysis of a 150-star subset from the SHINE survey, referred to as the F150 sample, Langlois et al. (2021) evaluated the survey’s detection capabilities. Utilizing SPHERE’s IRDIS H-band mode, the median detection performance achieved a $5\sigma$ contrasts of 11.8 mag at 200 mas and 13.1 mag at 800 mas. These results were obtained by applying standard post-processing techniques, including principal component analysis (PCA; Soummer et al., 2012; Amara and Quanz, 2012) and the template-based locally optimized combination of images (TLOCI; Marois et al., 2014).
@@
In recent years, more advanced post-processing algorithms such as ANDROMEDA (Cantalloube et al., 2015), PACO (Flasseur et al., 2018, 2020, 2024), SODINN / NA-SODINN (Gomez Gonzalez et al., 2018; Cantero et al., 2023), regime switching model (RSM; Dahlqvist et al., 2020, 2021b, 2021a), and 4S (Bonse et al., 2025) have been developed, with the goal of optimizing noise modeling and enhancing planetary signal detection – particularly for faint companions at close separations. Notably, Chomez et al. (2023) applied PACO to the SHINE F150 sample, improving the median contrast limits by a factor of five at inner working angles and a factor of two overall, and demonstrating the impact of more refined statistical modeling.
@@
Building on this progress, we revisit the SHINE F150 sample using the RSM algorithm, which has shown excellent performance in the Exoplanet Imaging Data Challenge (EIDC; Cantalloube et al., 2020b) by achieving high F1 scores and a low false positive rate. Unlike spatial-only modeling approaches, RSM captures both the spatial and temporal evolution of pixel intensities across residual, de-rotated frames using a Markov regime-switching framework applied to outputs from multiple PSF subtraction techniques. This enables it to fully exploit the temporal structure of high-contrast imaging sequences, offering improved sensitivity and robustness in exoplanet detection. In this study, we define detection thresholds tailored to RSM maps while accounting for noise variations across observing conditions. We compare two thresholding strategies, one assuming a lognormal noise distribution, and another based on maximizing the F1 score, and we assess their relative performance. We further evaluate how different post-processing algorithm combinations affect RSM sensitivity and present updated detection limits for the SHINE survey based on this reanalysis.
@@
The structure of this paper is as follows: Sect. 2 outlines the F150 sample selection criteria. In Sect. 3 we describe the observational conditions in the dataset. Section 4 presents an explanation of the RSM algorithm and its application in this study. In Sect. 5 we examine two methods for setting detection thresholds and their atmospheric context. In Sect. 6 we present contrast curves for different detection thresholds and post-processing combinations, highlighting the optimal approach within the RSM framework, and we discuss the improvements in the SHINE detection limits. In Sect. 7 we present the point sources detected by RSM in the F150 sample and compare them with those identified by other algorithms, such as PACO, and we include a discussion of newly detected candidates.
@@ 2 The data sample
In this study, we use the pre-reduced F150 sample from the SHINE survey, observed in H23 bands with ADI mode, covering an approximately $9\arcsec\times 9\arcsec$ field of view. The High Contrast Data Center (HC-DC, Delorme et al., 2017; Galicher et al., 2018; Beuzit et al., 2019; Maire et al., 2016) pre-processed these observations, resulting in a dataset of 343 observations. Atmospheric conditions during these observations varied significantly, with seeing values ranging from $0\aas@@fstack{\prime\prime}4$ to $3\arcsec$ and Strehl ratios from 0.1 to 0.95. To ensure data quality and consistency, a preselection of datasets was necessary based on these atmospheric parameters.
The preliminary selection of data was based on several parameters, including seeing, Strehl ratio, raw contrast, number of frames, full width at half maximum (FWHM), and an assessment of the features in the PSF and the science cubes. Information such as seeing, wind speed and coherence time were primarily derived from differential image motion monitor (DIMM, Sarazin and Roddier, 1990) measurements, as recommended by Milli et al. (2017), while the Strehl ratio was empirically measured from the PSF. Additional parameters, such as raw contrasts for science cubes and frame quality assessments, were provided by the HC-DC reduction pipeline. In cases where DIMM seeing measurements were unavailable, they were inferred from the available seeing values provided by the SPHERE real-time computer SPARTA (Fedrigo et al., 2006) using the relationship between DIMM and SPARTA seeing established in Milli et al. (2017).
To ensure a sample of stars with observing conditions ranging from fair to excellent, observations that met any of the following conditions were also discarded: seeing larger than $2\arcsec$, Strehl ratio less than 0.5, FWHM greater than 5.5 pixels, number of frames in the ADI sequence fewer than 40, and non-physical raw contrasts at 500 mas (negative or greater than one). In case of multiple observations of the same star within five consecutive nights, only the observation with the best conditions was retained. Finally, we assessed the quality of both coronagraphic and non-coronagraphic cubes using a combination of tools: the HC-DC frame quality assessment, correlation functions from the VIP package (Gomez Gonzalez et al., 2017; Christiaens et al., 2023), and visual inspection. Cubes containing a majority of corrupted frames were discarded. In addition, science cubes in which the star was located outside the coronagraph mask, or where the off-axis PSF showed strong secondary lobes caused by the low-wind effect (see Milli et al., 2018), were also excluded.
These filtering steps resulted in a total of 213 observations for 150 stars. These constitute the data sample used in this paper. Additional information about the list of targets and their observing conditions is provided in Section Data availability.
@@ 3 Clustering based on environmental conditions
The quality of observations and specific features within coronagraphic images can be characterized by several environmental and instrumental parameters. Factors such as seeing, wind speed, coherence time, Strehl ratio, and raw contrast collectively indicate observation quality and provide insights into SPHERE-specific features. For instance, the wind-driven halo (Cantalloube et al., 2020a) arises under high wind velocity and short coherence times, while the low-wind effect (LWE; Milli et al., 2018) appears at low wind speeds. These features can severely degrade SPHERE image quality, reducing the achievable contrast levels (see Cantalloube et al., 2019, for more details).
@@
We adopted the clustering method outlined in Dahlqvist et al. (2022), grouping observations with similar characteristics using the k-means clustering algorithm from the scikit-learn package. This approach not only helps identify distinct classes of SPHERE observations but also reduces computational load when optimizing PSF subtraction techniques and RSM parameters (as described in Section 4). The parameters selected for clustering include the number of frames in the ADI sequence, seeing, wind speed, parallactic angle range, Strehl ratio, raw contrast at 500 mas, and coherence time. Observations were grouped into six clusters. The number of clusters was determined by maximizing the silhouette score, a metric used to evaluate the quality of clustering. For each cluster, k-means automatically selected a representative observation, i.e., the dataset closest to the cluster centroid. In the following, the terms cluster center and representative observation are used interchangeably. The parameters for PSF subtraction and the RSM algorithm could then be optimized for these representative datasets, and generalized to other observations within the same cluster. To ensure unbiased parameter selection, we made sure that none of the cluster centers contained datasets with bright astrophysical signals. A pixel correlation test on science cubes was used to validate the effectiveness of clustering in capturing noise features (see Appendix A), and it supports the use of the k-means elected centers for parameter optimization. Figure 1 displays a $200\times 200$ pixel crop of representative individual frames from different clusters. These images illustrate the varying noise behavior across clusters, with some frames showing secondary lobes at close separations from the star indicative of the low-wind effect (Fig. 1a), while others exhibit strong elongated (i.e. butterfly) patterns characteristic of the wind driven halo (Fig. 1e).
@@
Figure 1: Images of single frames from various observations across different clusters. The images are labeled from (a) to (f), which correspond to clusters 1 to 6 presented in log scale.
@@
To visualize the clustering effectiveness based on selected parameters, we applied Principal Component Analysis (PCA) for dimensionality reduction (see Appendix B), enabling clear separation of clusters in the latent space. For a more physically grounded interpretation of the nature of the various clusters, Fig. 2 presents their visualization based on key environmental parameters, revealing distinct traits for each cluster:
@@
This successful clustering of the data into these meaningful categories allowed us to analyze noise distributions and detection limits based on well-defined observational conditions, and enabled a more targeted approach to interpreting our results.
@@ 4 The RSM algorithm
The RSM is a well-established econometric algorithm that was successfully adapted to HCI by Dahlqvist et al. (2020) to enhance the detection of faint companions at small angular separations from their host stars. In the RSM algorithm, patches of pixels from derotated, residual frames after PSF subtraction are used to construct time series that capture spatial and temporal noise behavior at a given separation. When a planetary signal is present, these time series exhibit significant deviations from the noise patterns across the pixel patch and instead follow a planetary model that is both spatially and temporally structured. The algorithm models the data with two regimes: a noise regime, describing the mean and variance of noise at a given separation, and a planetary regime, incorporating the noise profile to which is added a model of the PSF whose intensity, $\beta$, is tuned. Each central pixel of the patch in a given annulus is assigned a likelihood of belonging to either regime. The RSM algorithm employs a Markov Chain model, where the probability of a pixel belonging to the planetary regime depends on the likelihood of the pixel-centered patch alignment with a planetary model, the probability of adjacent pixels, and a transition probability that connects these points. This integration of spatial and temporal constraints, combined with the transition probabilities, enables the algorithm to reliably distinguish speckles from true astrophysical signals. Additionally, the RSM algorithm utilizes outputs from various PSF subtraction techniques to construct a unified time series for each annulus. By drawing on the temporal noise diversity across different post-processing methods, this approach enhances the sensitivity to faint companions.
@@
RSM has demonstrated significant advancements over current post-processing techniques, as highlighted in multiple studies (Dahlqvist et al., 2020, 2021b, 2022). Furthermore, RSM shows excellent performance compared to other post-processing algorithms, achieving the lowest false positive rate in the Exoplanet Imaging Data Challenge (EIDC; Cantalloube et al., 2020b), underscoring its reliability in distinguishing planetary signals from noise.
@@
One of the most sensitive aspects of the RSM algorithm involves selecting optimal parameters for both PSF subtraction techniques and RSM itself, including the planetary flux multiplicative factor $\delta$ (where planetary flux is expressed as a multiple of the standard deviation of pixel intensity at a given separation, $\beta=\delta\sigma$, see equation 1 in Dahlqvist et al., 2020), PSF model crop size, and the region used for noise estimation within an annulus. Dahlqvist et al. (2021a) introduced an automated optimization process within RSM to identify these optimal settings by maximizing contrast for PSF subtraction parameters and minimizing false positives for RSM parameters using the reversed parallactic angles, thereby enhancing the algorithm’s reliability in exoplanet detection. In this study, we employ RSM with a simplified parameter set, as recommended by Dahlqvist et al. (2022). In the following, we refer to the value of each pixel in the final RSM map – previously defined as the RSM probability – as the RSM score, to avoid any potential ambiguity.
@@
To reduce computational load and mitigate potential issues with the RSM parameter optimization on certain pathological datasets, we propose to determine the optimal parameters for the cluster centers identified in Sect. 3 and to apply these settings consistently across all datasets within each cluster. Table 1 in Appendix C details the specific parameters for each cluster. To assess the reliability of this approach, Appendix D compares the cluster center parameters with those computed individually for each dataset, providing insights into their alignment and the robustness of the clustering methodology. We exclude from the optimization process the selection of the optimal combination of PSF subtraction techniques within the auto-RSM framework, as delivered by the opti-combination method described in Dahlqvist et al. (2021a). Instead, we keep this parameter adjustable and explore how varying combinations of PSF subtraction techniques impact noise distributions and contrast curves across separations, as discussed in Sect. 6.